OpenStudy (jkbo):

Please help.solve the following system by graphing and identify the point of intersection. -5x + 2y = -4 2x - y = 2 A.(–2, 0) B.(1, 2) C.(2, 1) D.(0, –2)

OpenStudy (jdoe0001):

if you were to solve by "y" say -5x + 2y = -4 what would that give you?

OpenStudy (anonymous):

Do you mean solve "for" y?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

\[-5x +2y =-4\] \[2y =5x\] \[y =\frac{ 5 }{ 2 }x -2\] This would give you our beloved "favorite" form of a line; Slope-Intercept form OR y=mx+b

OpenStudy (anonymous):

It cut off my middle equation of 2y=5x-4

OpenStudy (jkbo):

thank you both

OpenStudy (jkbo):

but I m not sure if it is B. or C. @jdoe0001

OpenStudy (jdoe0001):

well, the "y" for the 1st one is correct as anthonyc123 showed y = 5/2 x -2 the 2nd one is y = 2x -2 now just graph those 2 the solution is where they cross each other

OpenStudy (jkbo):

I got D. is that correct?